OpenStudy (anonymous):

find the unit's digit in the product (2467)^153 * (341)^72

OpenStudy (phi):

341 to any power will always have a 1 in the unit's position (1*1=1)

OpenStudy (phi):

for 2467, concentrate on 7 to a power: 7^0 =0 7^1 = 7 7^2 = 49 7^3 = 343 7^4 = 2401 7^5 = 16807 looking at just the unit's digit we see it repeats after 7^4

OpenStudy (phi):

that means we should write 2467^153 as 2467^(4*38) * 2467^1 and then \[ \left(2467^4\right)^{38} \cdot 2467 \]

OpenStudy (phi):

as noted, the 2467^4 ends in 1, and any power of it ends in 1 we are left with a 7 from 2467 as the unit's digit of the product